Accounting Equation

Determine the missing amount for each of the following:

Assets = Liabilities +     Stockholders' Equity
a. $ = $229,650 + $717,660
b. $114,830 = $ + $35,880
c. $55,260 = $7,890 + $
Answer #1

Answer:

Assets = Liabilities + Stockholders Equity

a.Assets = 229650 + 717660 = 947310

b.Liabilities = Assets - Stockholders Equity = 114830 - 35880 = 78950

c.Stockholders Equity = Assets - Liabilities = 55260 - 7890 = 47370